OpenStudy (smarty):

Find the ordered pair where the maximum value occurs for the equation C = 12x – 4y given the following constraints:

OpenStudy (smarty):

\[y \ge4\] \[x \ge1\] \[x+2y\] \[y-x \le5\]

OpenStudy (smarty):

@zepp

OpenStudy (anonymous):

is it just x+2y?

OpenStudy (smarty):

\[x+2y \le16\]

OpenStudy (anonymous):

better...

OpenStudy (anonymous):

from that we get\[y \ge4,y \ge (12-C)/4,y \le 8-x/2, y \le 5+x\] intersecting, 8-x/2=5+x -> x=2,y=7 ,C=24-28=-4

OpenStudy (anonymous):

y=4, y=8-x/2 ,y=5+x

OpenStudy (anonymous):

intersecting 8-x/2=4 -> x=8, y=4 ,C=8*12-16=80

OpenStudy (anonymous):

intersecting 5+x=4 -> x=-1,y=4 -> C=-12-16=-28 Thus, answer is C=80

OpenStudy (anonymous):

Just to check if c=80 is valid, y>= (12-c)/4, c=80, y=4 'ok'

OpenStudy (smarty):

A.(1, 4) B.(2, 7) C.(8, 4) D.(10, 1)

OpenStudy (anonymous):

(2,7) -> c=80

OpenStudy (smarty):

hmk. Thanks.

OpenStudy (anonymous):

you plot those 3 eq, that will get u a triangle, find those vertices and plug the values into C, see which one has the higher value, double check if C is valid and that's it.

OpenStudy (smarty):

Alright, i'll do that on my grapghin calc

OpenStudy (anonymous):

you should get this triangle,